The length of a rectangle is 3 feet more than 4 times the width. If the area is 45 square feet, find the width and the length.

Answer by Cromlix(4381)

Width = x
Length = 3 + 4x
Area = width*length
45 = x(3 + 4x)
45 = 3x + 4x^2
4x^2 + 3x - 45 = 0
(4x + 15)(x - 3) = 0
4x + 15 = 0
x = -15/4 (no answer as x<0)
x - 3 = 0
x = 3
Width = 3 feet
Length = 15 feet
Hope this helps.
:-)